OpenStudy (anonymous):

Find the product. (n 3)2 · (n 5)4

OpenStudy (anonymous):

What is (n 3) 2...do you mean (n^3)^2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then, use the law of exponents, (a^x)^y = a^(xy)

OpenStudy (anonymous):

So what will be (n^3)^2?

OpenStudy (anonymous):

Find the product. (n 3)^2 · (n 5)^4

OpenStudy (anonymous):

Find the product. (n^ 3)^2 · (n ^5)^4

OpenStudy (anonymous):

As I mentioned before, the law of exponents states that (a^x)^y = a^(xy) Example: (4^3)^5 = 4^15 (you multiply exponents)

OpenStudy (anonymous):

(n^3)2(n^5)4 Multiply 3 by 2 to get 6. More Detail n^6(n^5)4 Multiply 5 by 4 to get 20. More Detail (n^6)⋅n^20 Multiply n^20 by each term inside the parentheses. More Detail n^26

OpenStudy (charlotte123):

you are right http://www.wolframalpha.com/input/?i=%28n%5E+3%29%5E2+%C2%B7+%28n+%5E5%29%5E4 :) @mathdummy23

OpenStudy (anonymous):

@charlotte123 thank you :)

OpenStudy (charlotte123):

no problem! good job helping! :D @mathdummy23

OpenStudy (jdoe0001):

\(\bf (n^3)^2\cdot (n^5)^4\implies n^{3\cdot 2}\cdot n^{5\cdot 4}\implies n^{3\cdot 2+5\cdot 4}\)